### v3

# Rate-In: Information-Driven Adaptive Dropout Rates for Improved Inference-Time Uncertainty Estimation

Rate-In: Information-Driven Adaptive Dropout Rates for Improved
Inference-Time Uncertainty Estimation
Tal Zeevi1, Ravid Shwartz-Ziv2, Yann LeCun2,3,
Lawrence H. Staib1†, John A. Onofrey1†
1Yale University, 2New York University, 3Meta FAIR
[…]
Accurate uncertainty estimation is crucial for deploying neural networks in risk-sensitive applications such as medical diagnosis. Monte Carlo Dropout is a widely used technique for approximating predictive uncertainty by performing stochastic forward passes with dropout during inference. However, using static dropout rates across all layers and inputs can lead to suboptimal uncertainty estimates, as it fails to adapt to the varying characteristics of individual inputs and network layers. Existing approaches optimize dropout rates during training using labeled data, resulting in fixed inference-time parameters that cannot adjust to new data distributions, compromising uncertainty estimates in Monte Carlo simulations.
In this paper, we propose Rate-In, an algorithm that dynamically adjusts dropout rates during inference by quantifying the information loss induced by dropout in each layer’s feature maps. By treating dropout as controlled noise injection and leveraging information-theoretic principles, Rate-In adapts dropout rates per layer and per input instance without requiring ground truth labels. By quantifying the functional information loss in feature maps, we adaptively tune dropout rates to maintain perceptual quality across diverse medical imaging tasks and architectural configurations. Our extensive empirical study on synthetic data and real-world medical imaging tasks demonstrates that Rate-In improves calibration and sharpens uncertainty estimates compared to fixed or heuristic dropout rates without compromising predictive performance. Rate-In offers a practical, unsupervised, inference-time approach to optimizing dropout for more reliable predictive uncertainty estimation in critical applications.
{NoHyper}footnotetext: These authors contributed equally to this work.
1 Introduction
Figure 1: Rate-In maintains anatomically meaningful uncertainty estimates at high dropout rates, while standard methods degrade into noise. Left: Prostate MRI segmentation showing ground truth (yellow), model predictions (blue), and errors (red). Right: Uncertainty maps (brighter regions indicate higher uncertainty) at increasing dropout rates (1%-16%). While constant dropout produces diffuse uncertainty that obscures anatomical details, Rate-In precisely highlights challenging regions, particularly along tissue boundaries, preserving clinical utility of the uncertainty estimates.
Deep neural networks have achieved remarkable success in a wide range of applications, from image recognition to natural language processing. However, their deployment in risk-sensitive domains such as medical diagnosis, autonomous driving, and financial forecasting requires not only accurate predictions but also reliable estimates of predictive uncertainty [1]. Accurate uncertainty estimation enables models to express confidence in their predictions, allowing practitioners to make informed decisions and to identify cases where the model may be less reliable.
Monte Carlo Dropout [2] is a popular and practical technique for approximating predictive uncertainty in neural networks. By performing stochastic forward passes with dropout activated during inference, Monte Carlo Dropout samples from an approximate posterior distribution over the network’s weights, enabling the estimation of both the mean prediction and the associated uncertainty. This method offers a balance between computational efficiency and uncertainty estimation quality, making it suitable for large-scale applications.
A common limitation of Monte Carlo Dropout is the use of fixed dropout rates during inference, typically the same rates used during training. These static dropout rates are applied uniformly across all layers and inputs, regardless of the varying characteristics of individual inputs and the differing sensitivities of network layers. This rigidity can lead to suboptimal uncertainty estimates [3]: fixed dropout rates may either inject excessive noise, degrading the model’s predictive performance, or fail to introduce sufficient variability to capture the true predictive uncertainty.
As illustrated in Figure 1, the way we estimate this uncertainty can dramatically affect our ability to identify truly challenging regions in tasks such as medical image segmentation. Traditional uncertainty estimation methods often produce overly diffuse uncertainty patterns that fail to distinguish between genuinely difficult decisions and simpler ones. This limitation becomes particularly critical in medical applications, where highlighting truly uncertain regions can guide clinical attention to areas requiring additional scrutiny.
Existing approaches have explored optimizing dropout rates during training [4, 5, 6], often relying on labeled data to learn optimal dropout parameters. However, these methods result in fixed dropout rates that remain static during inference, unable to adapt to new or diverse input distributions encountered after deployment. Moreover, in many practical scenarios, especially in medical applications, ground truth labels may not be available during inference, necessitating methods that can adapt without supervision.
To address these limitations, we propose leveraging information theory to dynamically adjust dropout rates during inference. By viewing dropout as controlled noise injection and each layer as a noisy communication channel [7], we can quantify the information loss induced by dropout in each layer’s feature maps. This perspective allows us to adjust dropout rates per layer and per input instance, aiming to maintain a desired level of information retention while introducing sufficient variability for uncertainty estimation.
In this paper, we introduce Rate-In, an algorithm that dynamically adjusts dropout rates during inference by quantifying the functional information loss in feature maps induced by dropout. Rate-In operates without requiring ground truth labels, making it suitable for deployment in real-world applications where labeled data is unavailable or scarce. By treating each layer’s dropout-induced information loss as a controllable parameter, Rate-In adaptively tunes dropout rates to maintain perceptual quality across diverse tasks and architectures.
Our main contributions are:
  • We introduce an information-theoretic framework for adjusting dropout rates during inference, interpreting dropout as controlled noise injection and quantifying its impact on information flow within the network.
  • We propose Rate-In, an algorithm that dynamically adjusts dropout rates per layer and per input instance by controlling the information loss induced by dropout, without requiring ground truth labels.
  • We demonstrate the effectiveness of Rate-In through extensive empirical evaluations on synthetic data and real-world medical imaging tasks, showing that it improves uncertainty estimation and calibration compared to fixed or heuristic dropout rates, without compromising predictive performance.
The rest of the paper is organized as follows: In Section 2, we provide background on uncertainty estimation in neural networks, discuss dropout and its limitations, and explore the application of information theory to deep learning. In Section 3, we detail the Rate-In algorithm and its implementation. Section 4 presents experimental results demonstrating the benefits of Rate-In in uncertainty estimation tasks. Finally, in Section 5, we discuss the implications of our findings and potential directions for future work.
2 Background
In this section, we provide the necessary background to understand our proposed method. We begin by discussing the importance of uncertainty estimation in neural networks and the existing approaches to achieve it. We then go over the Monte Carlo Dropout as a practical method for uncertainty estimation, highlighting its limitations due to fixed dropout rates during inference. Finally, we introduce the application of information theory in deep learning, particularly how it can inform the adaptive adjustment of dropout rates to improve uncertainty estimation.
2.1 Uncertainty Estimation in Neural Networks
Deep neural networks have revolutionized numerous fields due to their remarkable predictive capabilities. However, their deterministic nature often prevents them from expressing uncertainty about their predictions [8]. In risk-sensitive applications such as medical diagnosis, autonomous driving, and financial forecasting, understanding the confidence of a model’s predictions is as crucial as the predictions themselves [1]. Reliable uncertainty estimation allows practitioners to make informed decisions, identify cases where the model may be less reliable, and potentially defer to human experts when necessary.
Uncertainty in neural networks can be broadly categorized into two types [9]:
  • Aleatoric Uncertainty: Arises from inherent data noise and represents the uncertainty due to the stochastic nature of the data generation process. It is irreducible and cannot be eliminated even with more data.
  • Epistemic Uncertainty: Stems from model uncertainty, often due to limited data or model capacity, and can be reduced with more data or better modeling.
Accurately estimating both types of uncertainty is essential for building reliable models. Various methods have been proposed for uncertainty estimation in neural networks:
  • Bayesian Neural Networks (BNNs): Introduce probability distributions over network weights [10], allowing the model to capture uncertainty through posterior distributions. However, BNNs are often computationally expensive and challenging to scale.
  • Ensemble Methods: Train multiple models and aggregate their predictions [11]. Ensembles can capture both types of uncertainty but require significant computational resources.
  • Monte Carlo Dropout: Applies dropout at inference time to approximate Bayesian inference [2], providing a practical and computationally efficient method for uncertainty estimation.
2.2 Monte Carlo Dropout for Uncertainty Estimation
Monte Carlo (MC) Dropout leverages the dropout regularization technique [12] for uncertainty estimation by performing multiple stochastic forward passes through the network with dropout activated during inference [2]. This approach approximates a Bayesian posterior over the model parameters, allowing for the estimation of predictive uncertainty.
Given an input $\mathbf{x}$ and a neural network with parameters $\mathbf{W}$, the predictive distribution is approximated by averaging the predictions over $T$ stochastic forward passes:
$p(\mathbf{y}|\mathbf{x})\approx\frac{1}{T}\sum_{t=1}^{T}p(\mathbf{y}|\mathbf{x},\mathbf{W}_{t})$ (1)
where $\mathbf{W}_{t}$ represents the parameters of the network after applying dropout at iteration $t$. The variability in the predictions reflects the model’s uncertainty.
2.3 Limitations of Fixed Dropout Rates
While MC Dropout is a practical method for uncertainty estimation, it commonly uses fixed dropout rates across all layers and inputs during inference [2]. This approach has limitations:
  • Suboptimal Uncertainty Estimation: Fixed dropout rates may not adequately capture the predictive variance needed for accurate uncertainty estimation [3]. Some inputs may require higher dropout rates to reflect uncertainty, while others may suffer from excessive noise.
  • Potential Performance Degradation: Applying the same dropout rate uniformly can introduce unnecessary noise in some layers, potentially affecting model performance [12]. Critical features may be disrupted, adversely impacting prediction accuracy.
  • Lack of Adaptability: Fixed rates do not account for the varying complexity of different inputs or the differing sensitivities of layers within the network. This rigidity limits the model’s ability to adapt to new or diverse data distributions.
These limitations highlight the need for methods that can adapt dropout rates dynamically during inference to improve uncertainty estimation without compromising performance.
2.4 Information Theory in Deep Learning
Information theory provides a powerful framework for quantifying the amount of information transmitted through a system [7]. In the context of deep learning, information-theoretic concepts have been used to understand and improve neural networks in several ways:
  • Information Bottleneck Principle: Proposes that layers in a neural network aim to compress input data while preserving relevant information for the output [13]. This principle has been used to analyze learning dynamics and generalization.
  • Mutual Information Analysis: Estimating mutual information between inputs, hidden layers, and outputs helps in understanding how information is processed and transformed within the network [14, 15].
  • Regularization Techniques: Information-theoretic regularization methods, such as information dropout [16], have been introduced to improve robustness and generalization by controlling the flow of information during training.
However, most of these applications focus on the training phase of neural networks, using information theory to guide learning and improve representations. The potential of information theory to inform adaptive mechanisms during inference, particularly for uncertainty estimation, remains underexplored.
2.5 Applying Information Theory to Dropout
Interpreting dropout through the lens of information theory allows us to view it as introducing controlled noise into the network, effectively treating each layer as a noisy communication channel [7]. This perspective provides valuable insights into how dropout affects information flow within the network:
  • Dropout as Noise Injection: Dropout reduces mutual information between a layer’s input and output by introducing random perturbations [12, 16]. The noise can prevent overfitting by discouraging the network from relying too heavily on any single neuron.
  • Information Loss Quantification: The amount of information lost due to dropout can be quantified using measures like entropy and mutual information [17]. This quantification enables us to assess the impact of dropout on the network’s representations.
  • Layer Sensitivity: Different layers may exhibit varying sensitivities to information loss induced by dropout, depending on their role and the nature of their activations. Early layers may be more robust to noise, while later layers may be more sensitive.
These insights suggest that by controlling the information loss induced by dropout, we can adaptively adjust dropout rates to balance the trade-off between introducing variability for uncertainty estimation and maintaining sufficient information for accurate predictions.
3 Rate-In: Adaptive Dropout at Inference
In this section, we present Rate-In, an algorithm that dynamically adjusts dropout rates during inference by quantifying the information loss induced in each layer’s feature maps. By viewing dropout as controlled noise injection and leveraging information-theoretic principles, Rate-In adapts dropout rates per layer and per input instance, enhancing uncertainty estimation without compromising predictive performance.
3.1 Reinterpreting Dropout as Noise Injection
Traditionally, dropout serves as a regularization technique that randomly drops units or connections during training [12, 18]. In the context of uncertainty estimation during inference, we reinterpret dropout as a form of noise injection into the neural signal. This perspective allows us to analyze dropout’s impact on information flow within the network.
By considering each layer as a communication channel [7], applying dropout introduces noise into this channel, reducing the mutual information between the layer’s input and output. The extent of information loss depends on the dropout rate and the characteristics of the activations in that layer.
3.2 Quantifying Information Loss
To adjust dropout rates dynamically, we need to quantify the information loss induced by dropout in each layer. We employ the concept of mutual information (MI), which measures the shared information between two random variables [7].
Given a layer with input activations $\mathbf{h}{\text{in}}$ and output activations $\mathbf{h}{\text{out}}$, the mutual information $I(\mathbf{h}{\text{in}};\mathbf{h}{\text{out}})$ quantifies how much information about $\mathbf{h}{\text{in}}$ is preserved in $\mathbf{h}{\text{out}}$. Dropout reduces this mutual information due to the introduced randomness.
Directly computing MI in high-dimensional spaces is challenging. To address this, we use efficient estimation methods, such as the adaptive binning estimator [17], which approximates MI using entropy estimations with adaptive data binning.
3.3 The Rate-In Algorithm
Rate-In adjusts the dropout rate for each layer during inference by maintaining the information loss below a predefined threshold. The algorithm operates sequentially during the forward pass, adjusting dropout rates per layer and per input instance (Figure  2 and Algorithm  1).
Figure 2: Rate-In dynamically optimizes dropout rates through an information-theoretic feedback loop: (i) apply a dropout mask to the feature map; (ii) assess information integrity; (iii) calculate the difference from the target information level; (iv) update the dropout rate accordingly.
Algorithm 1 Rate-In Algorithm
1:Trained neural network $\mathcal{F}$ with $L$ layers, input $\mathbf{x}$, initial dropout rates $p^{(0)}$, information loss threshold $\epsilon$, maximum iterations $N_{\text{max}}$
2:Adjusted dropout rates $p$, output prediction $\hat{\mathbf{y}}$
3:Initialize $p\leftarrow p^{(0)}$
4:for $l=1$ to $L$ do
5:     $n\leftarrow 0$
6:     repeat
7:         Apply dropout with rate $p_{l}$ to layer $l$
8:         Compute activations $\mathbf{h}^{(l)}{\text{in}}$ and $\mathbf{h}^{(l)}{\text{out}}$
9:         Estimate mutual information $I(\mathbf{h}^{(l)}{\text{in}};\mathbf{h}^{(l)}{\text{out}})$
10:         Compute information loss $\Delta I_{l}=I^{(l)}{\text{full}}-I(\mathbf{h}^{(l)}{\text{in}};\mathbf{h}^{(l)}{\text{out}})$
11:         if $\Delta I_{l}>\epsilon$ then
12:              Decrease $p_{l}$
13:         else
14:              Increase $p_{l}$
15:         end if
16:         $n\leftarrow n+1$
17:     until $|\Delta I_{l}-\epsilon|<\delta$ or $n\geq N{\text{max}}$
18:end for
19:Perform forward pass with adjusted dropout to obtain $\hat{\mathbf{y}}$
20:return $p$, $\hat{\mathbf{y}}$
3.3.1 Algorithm Details
Initialization
We start with initial dropout rates $p^{(0)}$, which can be uniform across layers or based on prior knowledge. We define an information loss threshold $\epsilon$ specifying the maximum allowable mutual information loss per layer.
Layer-wise Adjustment
For each layer $l$, we perform the following steps:
  1. 1.
    Apply Dropout: Apply dropout with the current rate $p_{l}$ to the activations in layer $l$.
  2. 2.
    Compute Activations: Obtain the input and output activations $\mathbf{h}^{(l)}_{\text{in}}$ and $\mathbf{h}^{(l)}_{\text{out}}$ for layer $l$.
  3. 3.
    Estimate Mutual Information: Estimate the mutual information $I(\mathbf{h}^{(l)}_{\text{in}};\mathbf{h}^{(l)}_{\text{out}})$ using an appropriate estimator.
  4. 4.
    Compute Information Loss: Calculate the information loss $\Delta I_{l}=I^{(l)}_{\text{full}}-I(\mathbf{h}^{(l)}_{\text{in}};\mathbf{h}^{(l)}_{\text{out}})$, where $I^{(l)}_{\text{full}}$ is the mutual information without dropout.
  5. 5.
    Adjust Dropout Rate: If $\Delta I_{l}>\epsilon$, decrease $p_{l}$ to reduce information loss; otherwise, increase $p_{l}$ to introduce more variability.
  6. 6.
    Convergence Check: Repeat the adjustment until the information loss is within an acceptable range or the maximum number of iterations is reached.
Forward Pass
After adjusting dropout rates, perform a final forward pass with the adjusted rates to obtain $\hat{\mathbf{y}}$.
3.3.2 Implementation Considerations
Estimating Mutual Information
To make MI estimation tractable in high-dimensional spaces, we can:
  • Use dimensionality reduction to project activations to lower dimensions.
  • Employ non-parametric estimators like k-nearest neighbors or kernel density estimators.
  • Approximate MI using simpler metrics such as correlation or variance when appropriate.
Optimization Strategy
Adjusting dropout rates can be formulated as an optimization problem. Simple strategies like gradient descent or rule-based adjustments (e.g., fixed step sizes) can be used to reach the desired information loss threshold.
Computational Overhead
To mitigate the additional computations:
  • Limit the number of iterations per layer ($N_{\text{max}}$).
  • Share computations across similar inputs or batches.
  • Precompute or approximate certain quantities.
3.4 Discussion of the Rate-In Approach
The Rate-In algorithm offers several advantages:
  • Adaptive Dropout Rates: By adjusting dropout rates per layer and per input instance, Rate-In accommodates the varying sensitivities of layers and the complexities of different inputs.
  • Improved Uncertainty Estimation: By controlling the information loss, Rate-In ensures that sufficient variability is introduced for uncertainty estimation without degrading the predictive performance.
  • No Need for Ground Truth Labels: Rate-In operates during inference without requiring labeled data, making it applicable in deployment scenarios where labels are unavailable.
However, there are also challenges:
  • Computational Complexity: The need to estimate mutual information and adjust dropout rates can increase inference time.
  • Estimation Accuracy: Accurate estimation of mutual information is non-trivial, especially in high-dimensional spaces.
  • Hyperparameter Selection: Choosing appropriate values for the information loss threshold $\epsilon$ and other hyperparameters requires careful consideration.
In the next section, we present experimental results demonstrating Rate-In’s effectiveness across various tasks and architectures.
4 Experiments
In this section, we evaluate the effectiveness of the proposed Rate-In algorithm for predictive uncertainty estimation through MC simulations. We conduct experiments on both synthetic data and real-world medical imaging tasks, utilizing state-of-the-art pre-trained neural network architectures. Our evaluation focuses on three main objectives:
  1. (i)
    Effectiveness of Rate-In: Assess Rate-In’s ability to improve uncertainty estimation and calibration compared to fixed or heuristic dropout rates.
  2. (ii)
    Impact on Predictive Performance: Determine whether dynamically adjusting dropout rates at inference time affects predictive accuracy.
  3. (iii)
    Adaptability across Tasks and Architectures: Demonstrate the applicability of Rate-In across different tasks and neural network architectures.
4.1 Experimental Setup
We evaluate our method on three categories of tasks:
Synthetic Regression
To validate our method in a controlled setting, we generate a one-dimensional regression dataset with varying levels of noise. The target variable is defined as $y=\sin(x)+\epsilon$, where $x$ is uniformly sampled from $[-3,3]$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. We create datasets with five different noise levels by setting $\sigma\in{0.1,0.2,0.3,0.4,0.5}$. Each dataset contains 100 training and test instances.
Medical Image Classification
For real-world classification tasks, we utilize three datasets from the MedMNIST v2 repository [19]:
  • PathMNIST: 107,180 histopathology images for colorectal cancer tissue classification (9 classes).
  • BloodMNIST: 17,092 blood cell microscope images for peripheral blood cell classification (8 classes).
  • TissueMNIST: 236,386 kidney cortex microscope images for kidney cortex cell type classification (8 classes).
Medical Image Segmentation
We select two segmentation tasks involving different imaging modalities:
  • Prostate MRI Segmentation[20]: Segmenting the peripheral and transitional zones in prostate MRI scans.
  • Liver Tumor CT Segmentation[20]: Segmenting liver and tumor regions in CT scans.
Baselines Methods
We compare Rate-In with the following methods:
  • Fixed Dropout[12]: Applying a constant dropout rate across all layers during MC iterations.
  • Scheduled Dropout[21]: Linearly decreasing the dropout rate over MC iterations.
  • Activation-Based Dropout: Adjusting dropout rates based on the coefficient of variation (CoV) of layer activations, inspired by adaptive dropout methods [4].
4.1.1 Implementation Details
To directly compare Rate-In with classical dropout approaches, we set the information loss threshold $\epsilon$ equal to the fixed dropout rate $p^{(0)}$. $\delta$ was set to 0.01 based on preliminary experiments. The maximum number of iterations per layer was set to $N_{\text{max}}=30$.
For information loss estimation, we use the adaptive binning estimator [17], which efficiently approximates mutual information in high-dimensional feature spaces. Additionally, we measured the information using structural similarity index measure (SSIM) [22] (See Appendix B for more information and results). Monte Carlo simulation involves 30 forward passes per input, with dropout rates dynamically adjusted according to the measured information loss. We employ the following neural network architectures:
  • Synthetic Regression: A fully connected network with two hidden layers of 50 units each and ReLU activations.
  • Classification Tasks: ResNet-18 [23] pre-trained on MedMNIST datasets, with dropout layers added after each residual block’s ReLU activation.
  • Segmentation Tasks: nnU-Net [24], a U-Net-based architecture with dropout layers after each encoder and decoder block.
Evaluation Metrics
For predictive performance, we use Mean Squared Error (MSE) for regression, Accuracy (ACC) for classification, and Dice Similarity Coefficient (DSC) for segmentation. To evaluate uncertainty estimation, we employ Expected Calibration Error (ECE)[25], Area Under the Accuracy-Rejection Curve (AUARC)[26], and Boundary Uncertainty Coverage (BUC) [27]. We also provide qualitative analyses through uncertainty maps and reliability diagrams.
Our code and complete implementation details are available in the supplementary material. The appendix includes full training protocols, hyperparameter settings, architectural details, and ablation studies.
4.2 Results
4.2.1 Synthetic Data Analysis
We first evaluate Rate-In on the synthetic regression task to validate our approach under controlled conditions. Figure  3(a) shows the efficiency of uncertainty estimation across increasing noise levels ($\sigma\in\{0.1,0.2,0.3,0.4,0.5\}$) with a fixed training set size ($N=100$). The efficiency is quantified using the ratio of interval width to empirical coverage; lower ratios indicate more precise uncertainty estimates without sacrificing reliability. Rate-In consistently maintains a lower ratio across all noise levels compared to baseline methods.
Figure  3(b) illustrates the impact of training set size under a fixed noise level ($\sigma=0.01$). Rate-In outperforms other methods across all training sample sizes, achieving more efficient uncertainty estimation.The results demonstrate Rate-In’s ability to adapt dropout rates based on local uncertainty characteristics, leading to more efficient and reliable uncertainty estimation compared to static dropout approaches. The method shows particular advantages in scenarios with heteroscedastic noise and limited training data.
(a) Increasing noise level
(b) Increasing training points
Figure 3: Rate-In is more efficient in uncertainty estimation. The ratio of uncertainty interval width to 95% coverage. Lower ratios indicate more efficient uncertainty estimation. (a) Varying noise levels; (b) Varying number of training points at fixed noise level $(\sigma=0.01)$.
4.3 Classification Tasks
Table  1 summarizes the performance metrics for each classification task and dropout policy. Rate-In consistently achieves higher accuracy and better uncertainty calibration compared to baseline methods. Specifically, for the PathMNIST dataset, Rate-In maintains high performance across different levels of information loss, as illustrated in Figure 4. At a 0.05 dropout rate, Rate-In maintains full model accuracy for both TissueMNIST and PathMNIST datasets, while other methods show significant degradation.
The three medical datasets exhibit distinct patterns of robustness to dropout. BloodMNIST retains high performance even under significant dropout, while TissueMNIST shows greater sensitivity to information loss, and PathMNIST demonstrates intermediate resilience. These differences highlight the importance of adaptive dropout strategies across different medical imaging applications.
Table 1: Rate-In maintains near-original model accuracy even at high dropout rates, while baseline methods show severe degradation. Classification performance across three medical imaging datasets, reported as Accuracy (ACC) and AUARC. Results demonstrate that Rate-In consistently outperforms conventional dropout strategies, particularly at higher dropout rates. Higher AUARC values (in parentheses) indicate better uncertainty estimation quality.
Dataset Method ACC Full Model Dropout Rate
0.05 0.10 0.20
TissueMNIST Constant .68 .53 (.67) .31 (.37) .09 (.25)
Scheduled .56 (.70) .36 (.47) .13 (.19)
Activation .54 (.67) .32 (.40) .10 (.24)
Rate-In .68 (.84) .64 (.80) .48 (.59)
PathMNIST Constant .90 .87 (.96) .75 (.91) .41 (.56)
Scheduled .88 (.97) .78 (.93) .47 (.65)
Activation .87 (.96) .76 (.92) .44 (.60)
Rate-In .90 (.98) .88 (.97) .75 (.91)
BloodMNIST Constant .97 .95 (.99) .91 (.98) .58 (.75)
Scheduled .95 (.99) .92 (.98) .75 (.86)
Activation .96 (.99) .93 (.99) .81 (.92)
Rate-In .96 (.99) .93 (.99) .82 (.92)
Figure 4: Rate-In maintains high classification accuracy even at aggressive dropout rates where conventional methods degrade. Classification accuracy on the PathMNIST dataset comparing dropout strategies across increasing rates. Traditional methods show sharp decline, while Rate-In preserves performance by adaptively adjusting dropout patterns based on local feature importance.
4.4 Segmentation Tasks
Table 2: Rate-In achieves superior uncertainty estimation across all anatomical zones while preserving or improving segmentation accuracy compared to best benchmarks. Comprehensive evaluation across different anatomical regions in MRI and CT modalities, showing DSC, ECE, and BUC at varying dropout rates. Rate-In demonstrates consistent improvements in uncertainty metrics (lower ECE, higher BUC) while maintaining competitive DSC scores, with particularly strong performance in challenging regions like the transitional zone and tumor boundaries.
Zone Modality Method Full Model DSC (Full Model, % change) ECE ×$\times 10^{-3}$ BUC
DSC p=0.01 p=0.05 p=0.10 p=0.01 p=0.05 p=0.10 p=0.01 p=0.05 p=0.10
Peripheral MRI Rate-In 0.682 +0.29% +2.79% +2.35% 4.60 4.26 4.58 0.67 0.63 0.61
Best Benchmark +1.17% +2.79% +2.64% 4.62 4.40 5.01 0.58 0.53 0.50
Transitional MRI Rate-In 0.892 -0.34% -0.78% -0.67% 3.97 4.03 4.86 0.87 0.84 0.80
Best Benchmark +0.11% -0.78% -0.78% 4.25 6.60 10.73 0.70 0.62 0.57
Liver CT Rate-In 0.955 +1.05% +1.05% +0.94% 6.73 5.54 4.65 0.93 0.93 0.92
Best Benchmark 0.00% 0.00% 0.00% 6.90 5.80 5.00 0.92 0.91 0.91
Tumor CT Rate-In 0.579 +1.21% -0.52% -2.59% 1.78 1.78 1.88 0.61 0.53 0.48
Best Benchmark +0.17% -1.90% -1.73% 1.80 1.80 1.90 0.60 0.55 0.53
Rate-In’s advantages are particularly evident in medical image segmentation, where accurate uncertainty estimation is crucial. As illustrated in Figure 1, Rate-In generates uncertainty maps that precisely highlight anatomically challenging regions, especially along tissue boundaries where expert radiologists often face ambiguity. In contrast, constant dropout produces diffuse uncertainty patterns at higher dropout rates, obscuring areas of genuine uncertainty.
Rate-In demonstrates several key advantages:
  • Anatomical Precision: Rate-In accurately pinpoints ambiguous regions, particularly along tissue boundaries, enhancing clinical interpretability. Constant dropout yields diffuse patterns that obscure genuine concerns.
  • Dropout Rate Stability: Rate-In maintains consistent, focused uncertainty estimates across varying dropout rates, unlike conventional methods that degrade with increased rates, ensuring reliable clinical performance.
  • Error Correlation: Rate-In’s high-uncertainty regions strongly correlate with actual prediction errors, indicating better calibration and enhancing trust in the model’s predictions.
  • Clinical Interpretability: Rate-In’s uncertainty maps align with anatomical structures and boundaries, making them more interpretable. It facilitates the integration of statistical uncertainty into clinical decision-making.
These qualitative observations are supported quantitatively by performance metrics across our test set. Table 2 summarizes the metrics for each segmentation task and dropout policy. Rate-In consistently outperforms benchmarks in uncertainty metrics such as ECE and BUC across most anatomical zones. For instance, in regions with inherent boundary ambiguity, Rate-In shows enhanced capability in quantifying uncertainty at region interfaces, evidenced by higher BUC scores in the Transitional zone. This improved boundary characterization is critical for clinical applications where precise delineation between overlapping anatomical zones impacts diagnostic accuracy.
The results demonstrate that Rate-In’s information-theoretic approach to adjusting dropout rates produces uncertainty estimates that are both more accurate and more clinically meaningful. The method’s ability to maintain focused uncertainty patterns while preserving anatomical detail makes it particularly valuable for medical imaging applications where precision and reliability are paramount.
Figure 5 illustrates the impact of dropout rates on information loss across network layers. In the pre-trained U-Net model, the decoder layers experience a significant decline in information, with a reduction of approximately 40% as dropout rates increase from 5% to 20%. In contrast, the encoder layers are less affected, showing smaller changes in structural similarity with varying dropout rates. This suggests that Rate-In effectively allocates dropout rates to minimize information loss where it is most critical.
Figure 5: Decoder layers are more sensitive to information loss from dropout than encoder layers, demonstrating the need for layer-specific dropout rates. Analysis of information retention across U-Net layers as dropout rates increase. Decoder layers show significant information loss, while encoder layers maintain higher information. This asymmetric behavior validates Rate-In’s layer-adaptive approach, which can allocate lower dropout rates to sensitive decoder layers while maintaining higher rates in more robust encoder layers for efficient uncertainty estimation.
5 Summary
We introduced Rate-In, a novel algorithm that dynamically adjusts dropout rates during inference to enhance predictive uncertainty estimation in neural networks. Traditional methods like Monte Carlo Dropout use fixed dropout rates, which fail to adapt to individual inputs and network layers, leading to suboptimal uncertainty estimates. By treating dropout as controlled noise injection and leveraging information-theoretic principles, Rate-In quantifies the information loss in each layer and adjusts dropout rates accordingly, without requiring additional training or labels.
Our extensive experiments on synthetic data and real-world medical imaging tasks demonstrate that Rate-In consistently improves uncertainty calibration and sharpens uncertainty estimates compared to fixed or heuristic dropout rates, all while maintaining predictive performance. In segmentation tasks, Rate-In produced uncertainty maps that accurately highlighted anatomically challenging regions, enhancing clinical interpretability. By effectively minimizing critical information loss, Rate-In offers a practical, unsupervised inference-time solution for optimizing dropout, making it particularly valuable for risk-sensitive applications where accurate uncertainty estimation is essential.